NASA is conducting an experiment to find out the fraction of people who black out at G forces greater than 6. In an earlier study, the population proportion was estimated to be 0.3. How large a sample would be required in order to estimate the fraction of people who black out at 6 or more Gs at the 80% confidence level with an error of at most 0.03? Round your answer up to the next integer.

A= 383
Answer #1

Solution :

Given that,

\hat p= 0.3

1 - \hat p = 1 - 0.3 = 0.7

margin of error = E = 0.03

At 80% confidence level the z is ,

\alpha = 1 - 80% = 1 - 0.80 = 0.20

\alpha / 2 = 0.20 / 2 = 0.10

Z\alpha/2 = Z0.10= 1.28

Sample size = n = (Z\alpha/2 / E)2 * \hat p * (1 - \hat p )

= (1.28 / 0.03)2 * 0.3 * 0.7

= 383

Sample size = 383